OpenStudy (anonymous):

Find the partial derivative of the function f(x,y)= ((3x-2y)/(4x-4y))

OpenStudy (anonymous):

\[\frac{\partial f(x,y)}{\partial x}=-\frac{y}{4(x-y)^2}\] \[\frac{\partial f(x,y)}{\partial y}=-\frac{x}{4(x-y)^2}\]

OpenStudy (anonymous):

I tried putting those answers in but it says the answer isn't right.

OpenStudy (anonymous):

it wants me to find fx(x,y) and fy(x,y)

OpenStudy (zarkon):

the first one is correct...the second is not

OpenStudy (anonymous):

can you tell me how he got those answers?

OpenStudy (zarkon):

quotient rule

OpenStudy (anonymous):

ohhhhh i get it now. Thanks guys